OpenStudy (anonymous):

Suppose you own 60 shares of Microsoft, which pay a dividend of $0.13 per share per year. How much will you receive in dividends over 5 years, assuming the dividends stay the same and you buy no more stock? A. $39.00 B. $23.40 C. $44.60 D. $7.80

OpenStudy (anonymous):

=0.13*5=39

OpenStudy (anonymous):

The answer is correct, but the explanation should be .13*60*5 = 39.00